PLEASE HELP SOLVE
FACTOR EACH POLYNOMIAL
1) (H+K)^2-9
2) X^2+10X+25

Answer by venugopalramana(3286)

PLEASE HELP SOLVE
FACTOR EACH POLYNOMIAL
1) (H+K)^2-9
use formula....a^2-b^2=(a+b)(a-b)
(h+k)^2-3^2
=(h+k+3)(h+k-3)

2) X^2+10X+25...
x^2+5x+5x+25
x(x+5)+5(x+5)
(x+5)(x+5)=(x+5)^2